OpenStudy (anonymous):

Which equation represents a true proportion? Answer 5 over 8 = 9 over 12 5 over 8 = 9 over 16 3 over 4 = 9 over 12 3 over 4 = 9 over 16

OpenStudy (anonymous):

a proportion is a ratio where the top and bottom changes by the same factor, simply put: 3/4 = 9/12 because if you take 3/4 and you multiply the top by three and the bottom by 3, you get 9/12

OpenStudy (anonymous):

the third one.

OpenStudy (anonymous):

third one: 3/4 = 9/12